# An Experimental Evaluation of Imputation Models for Spatial-Temporal Traffic Data

### v1

@@ VI-E Case study
For the spatial random missing pattern, we randomly select a node to visualize the impute results and ground truth. As shown in Figure 10 (a) and (b), we observe that the overall imputation curve of each model is aligned with the ground truth, demonstrating their ability to capture the temporal dependence. Among them, BRITS, GCASTN, ImputeFormer, and LATC are more closely to real data. A common characteristic of these models is their incorporation of prior knowledge about missing data, which enhances imputation accuracy. BRITS and GCASTN leverage time-delay strategies to model temporal dependencies, while LATC and ImputeFormer adopt low-rank approaches to better capture the underlying data structure.
For the spatial continuous missing patterns, we select a cluster of missing nodes, visualize their connection relationships and observe the MAE metric of each node. The results as shown in Figure 10 (c) and (d). At SCTR missing pattern, ASTGNN, GCASTN, BRITS, and PriSTI have smaller errors, indicating the importance of modeling temporal dependencies, especially when spatial information is unavailable. We also notice that IGNNK performs poorly under the SCTC missing pattern due to its only focus on local spatial correlations and ignoring temporal dependencies, making it ineffective for this type of missing pattern.
VI-F Summary of New Insights
@@
VI-F1 Effective Components
(1) Incorporating prior knowledge to initialize missing values is important. To initialize missing values, BRITS and GCASTN employ the time delay mechanism, and PriSTI adopts the linear imputation mechanism. And they achieve more accurate imputation performance. Compared to them, IGNNK and mTAN use 0-initiation, exhibiting unsatisfactory performance. (2) Tensor completion, which captures global patterns like periodicity in traffic data, is useful. We can see that TC based ImputeFormer and LATC outperform most deep learning models in many scenarios.
VI-F2 Model Choosing in Various Scenarios
(1) Different missing patterns. GCASTN is the best model for temporal continuous missing pattern, and BRITS is the best model for spatial continuous missing pattern. (2) Memory-constrained scenarios. TC is the best choice since LATC has the smallest memory cost. Among deep learning methods, the diffusion model is the most appropriate choice for balancing performance and memory. (3) Scenarios with real-time requirements for inference speed. TC has the fastest inference speed. For deep learning methods, E2GAN is the fastest, and BRITS is the second.
VI-F3 Relations between Prediction and Imputation
@@
VII Conclusion
In this paper, we propose a practice-oriented categorization of imputation models. We classify the real traffic missing scenarios into four missing patterns and five missing rates. The imputation models are further organized based on their spatial-temporal modeling techniques and loss function designs. Additionally, we propose a unified pipeline to evaluate these models across three standardized traffic datasets and comprehensive missing patterns and analyze their performance in terms of efficiency, effectiveness, and different time intervals. Through extensive experimental analysis, we draw some valuable insights among effective components, model choosing in various scenarios, and the relation between imputation and prediction, which provide valuable guidance for future research.

### v2

@@ V-E Case study
For the spatial random missing pattern, we randomly select a node to visualize the impute results and ground truth. As shown in Figure 22(a) (a) and (b), we observe that the overall imputation curve of each model is aligned with the ground truth, demonstrating their ability to capture the temporal dependence. Among them, BRITS, GCASTN, ImputeFormer, and LATC are more closely to real data. A common characteristic of these models is their incorporation of prior knowledge about missing data, which enhances imputation accuracy. BRITS and GCASTN leverage time-delay strategies to model temporal dependencies, while LATC and ImputeFormer adopt low-rank approaches to better capture the underlying data structure.
For the spatial continuous missing patterns, we select a cluster of missing nodes, visualize their connection relationships and observe the MAE metric of each node. The results as shown in Figure 22(a) (c) and (d). At SCTR missing pattern, ASTGNN, GCASTN, BRITS, and PriSTI have smaller errors, indicating the importance of modeling temporal dependencies, especially when spatial information is unavailable. We also notice that IGNNK performs poorly under the SCTC missing pattern due to its only focus on local spatial correlations and ignoring temporal dependencies, making it ineffective for this type of missing pattern.
V-F Summary of New Insights
@@
V-F1 Suggestions for Model Design
(1) Incorporating prior knowledge is usually beneficial. There are two common strategies for integrating prior knowledge into imputation models. The first is to use prior knowledge to guide the imputation of missing values. For example, BRITS and GCASTN adopt a time-delay mechanism, while PriSTI first applies a linear imputation strategy. The second is to incorporate prior knowledge as a regularization term. For instance, LATC leverages a low-rank prior assumption and employs truncated nuclear normalization. Experimental results demonstrate that these models deliver strong and competitive performance. (2) Capture the global patterns such as periodicity, in traffic data is important. We observe that tensor completion–based models, such as ImputeFormer and LATC, which exploit these global patterns, outperform most other models across most scenarios.
V-F2 Suggestions for Model Selection
(1) Considering different missing patterns, GCASTN is the best model for temporal continuous missing pattern, and BRITS is the best model for spatial continuous missing pattern. (2) Considering memory-constrained scenarios, TC is the best choice since LATC has the smallest memory cost. Among deep learning methods, PriSTI, ImputeFormer, and IGNNK strike a good balance between performance and memory efficiency. (3) Considering the real-time requirements for inference speed. TC has the fastest inference speed. For deep learning methods, E2GAN is the fastest, and BRITS is the second fastest.
V-F3 Relations between Prediction and Imputation
@@
VI Conclusion
In this paper, we propose a practice-oriented taxonomy of imputation models. Specifically, the models are systematically categorized based on their spatial–temporal modeling techniques and loss function designs. Furthermore, we develop a unified benchmarking pipeline to evaluate these models on four standardized traffic datasets under diverse missing patterns. Their performance is comprehensively assessed in terms of efficiency, effectiveness, and robustness across during challenging periods with significant traffic fluctuations.
Through extensive experiments, we derive several practical insights regarding model design, model selection under different scenarios, and the relation between imputation and prediction, which can provide valuable guidance for future research and real-world deployment.